OpenStudy (dannyrod2000):

A recipe calls for 5 cups of flour and 3 cups of sugar. You want to make the recipe with 3.5 cups of flour. How much sugar should you use?

OpenStudy (dannyrod2000):

Step 1: Set up proportion. Step 2: Cross multiply and solve for the variable. Simplify. Step 3: Write the answer in a complete sentence.

OpenStudy (dannyrod2000):

ok so i got all of it exept Step 3

OpenStudy (dannyrod2000):

*except

OpenStudy (skullpatrol):

Write the answer in a complete sentence.

OpenStudy (dannyrod2000):

that is what i need help in.

OpenStudy (dannyrod2000):

i already have 5/3=3.5/x and 5x/5=10.5/5 and x=2.1

OpenStudy (dannyrod2000):

can you help me with that because i have no information about that.

OpenStudy (skullpatrol):

The question asks for: "How much sugar should you use?" Write an answer, to what the question asks for.

OpenStudy (dannyrod2000):

i need to use 2.1 cup of sugars.

OpenStudy (dannyrod2000):

like that?

OpenStudy (skullpatrol):

yes

OpenStudy (dannyrod2000):

ok thanks

OpenStudy (skullpatrol):

thanks for trying to learn :)

OpenStudy (dannyrod2000):

Thanks. :)

OpenStudy (dannyrod2000):

It just came to my mind. Is that your actual face?

OpenStudy (dannyrod2000):

It looks like a guy in a commercial.

OpenStudy (skullpatrol):

nah it's some model :)

OpenStudy (dannyrod2000):

o lol :)